please solve fast with correct answers.
Thank you.

QUESTION 15 Find the value of c+d if the angle between the vectors uz OA.2 O B.-1 OC.3 D. E.
Answer #1


1Page: 2 & 18 +18 d² + 2 + 36d-12cd-12c. = 54 + 202 + 27 d + c²d² 36d-12cd-12C = 36+ 9 d² + c²d² take doo – 126 = 36 = = -3 72